OpenStudy (mackenzie2013):

Presley has 6 books. There is enough space on a shelf for 4 books. In how many ways can 4 of the 6 books be arranged on the shelf?

jimthompson5910 (jim_thompson5910):

same idea as last time just with 4 slots this time (A,B,C,D) 6 choices for slot A 5 for B 4 for C 3 for D 6*5*4*3 = 30*12 = 360 different arrangements order matters in this case

OpenStudy (mackenzie2013):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome